OpenStudy (anonymous):

Simplify: (5x^8)(5x^7)/15x^4

OpenStudy (anonymous):

\[5x ^{19}/3\\]

OpenStudy (anonymous):

thanks :) can you help me with another one?

OpenStudy (anonymous):

Sure (;

OpenStudy (anonymous):

\[(\frac{6y^7 }{2y^2 })^{-4}\]

OpenStudy (anonymous):

Write In Normal

OpenStudy (anonymous):

(6y^7/2y^2)^-4

OpenStudy (anonymous):

\[\frac{ 1 }{ 81y ^{36} }\]

OpenStudy (anonymous):

its not the right answer :(

OpenStudy (anonymous):

Thats aa 36 BTW

OpenStudy (anonymous):

i know. i put it in but its the wrong answer

OpenStudy (anonymous):

Are You Sure You Told Me The Correct Problem ? Did You Leave Something Out?

OpenStudy (anonymous):

i didnt leave anything out. its just really confusing.

OpenStudy (mertsj):

\[(\frac{6y^7}{2y^2})^{-4}=(\frac{2y^2}{6y^7})^4=(\frac{1}{3y^5})^4=\frac{1}{81y ^{20}}\]